OpenStudy (anonymous):

what function would be the inverse of f(x) = The quantity of four x minus three, over two.?

OpenStudy (luigi0210):

Switch y and x, the solve for y.

OpenStudy (ivettef365):

so you have f(x) = 4x-3 ------ 2

OpenStudy (anonymous):

@ivettef365 yes

OpenStudy (ivettef365):

so the inverse you need to put it like this 4y-3 x = ------ solve for y 2

OpenStudy (ivettef365):

so you have 2x = 4y -3 2x + 3 = 4y 2x+3 so y = -------- 4

OpenStudy (anonymous):

@ivettef365 thats what i got i was just unsure lol thank you

OpenStudy (ivettef365):

your welcome :)